OpenStudy (anonymous):

Write an expression to describe a rule for the sequence. Then find the 100th term in the sequence. 5, 13, 21, 29, 37, 45, … (1 point) • 8n; 800 • 8n – 3; 797 • 3n – 8; 292 • n + 8; 108 my answer is C check my work plz

rishavraj (rishavraj):

\[a_n = a_1 + (n - 1)d \] \[a_{100} = 5 + (99)8 \]

OpenStudy (anonymous):

i need help, please look at my profile question

OpenStudy (anonymous):

so D?

rishavraj (rishavraj):

how it could be D... just solve the equation........ @farmgirl11

rishavraj (rishavraj):

\[a_100 = 5 + 99\times 8\] \[a_100 = 797\]

OpenStudy (anonymous):

so B